Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to Rs 400 and Rs 120.


1
Expert's answer
2020-11-17T07:19:09-0500

At equilibrium price, the quantity demanded is equals to the quantity supplied to the market.


This implies that Demand=Supply,


Qd=QsQd=Qs




1200-P = 120+3P1200−P=120+3P


Solving the above


4P = 10804P=1080


Hence Equilibrium price


P = Rs 270P=Rs270




When price Rises to Rs 400


Qs = 120 + (3 \times400)Qs=120+(3×400)


Qs = 1320Qs=1320


and


Qd = 1200 - 400 = 800Qd=1200−400=800


From the above price Rd 400, we can say that There is more supply than Demand because the price is high.




When Price rises to Rs 120


Qs = 120 + (3 \times120)Qs=120+(3×120)


Qs = Rs 480Qs=Rs480


And


Qd = 1200 - 120 = Rs 1080Qd=1200−120=Rs1080




The above price change implies that there is more demand than supply as the price is low.